Find the standard deviation for a security that has three one-year returns of -5%, 15%, and 20%.

5.00%

11.41%

13.23%

6.25%
Answer #1

Average return=Total return/Total time period

=(-5+15+20)/3=10%

Return (Return-Average return)^2
-5 (-5-10)^2=225
15 (15-10)^2=25
20 (20-10)^2=100
Total=350%

Standard deviation=[Total (Return-Average return)^2/(Time period-1)]^(1/2)

=[350/(3-1)]^(1/2)

=[350/2]^(1/2)

=13.23%(Approx).